# DepthSplat: Connecting Gaussian Splatting and Depth

### v2

@@ Abstract
Gaussian splatting and single/multi-view depth estimation are typically studied in isolation. In this paper, we present DepthSplat to connect Gaussian splatting and depth estimation and study their interactions. More specifically, we first contribute a robust multi-view depth model by leveraging pre-trained monocular depth features, leading to high-quality feed-forward 3D Gaussian splatting reconstructions. We also show that Gaussian splatting can serve as an unsupervised pre-training objective for learning powerful depth models from large-scale unlabeled datasets. We validate the synergy between Gaussian splatting and depth estimation through extensive ablation and cross-task transfer experiments. Our DepthSplat achieves state-of-the-art performance on ScanNet, RealEstate10K and DL3DV datasets in terms of both depth estimation and novel view synthesis, demonstrating the mutual benefits of connecting both tasks.
@@ 1 Introduction
The integration of 3DGS with single/multi-view depth estimation presents a compelling solution to overcome the individual limitations of each technique while at the same time enhancing their strengths. To this end, we introduce DepthSplat, which exploits the complementary nature of sparse-view feed-forward 3DGS and robust monocular depth estimation to improve the performance for both tasks.
@@
Thanks to our improved multi-view depth model, the quality of novel view synthesis with Gaussian splatting is also significantly enhanced (see Fig. 1 left). In addition, our Gaussian splatting module is fully differentiable, which requires only photometric supervision to optimize all model components. This provides a new, unsupervised way to pre-train depth prediction models on large-scale unlabeled datasets without requiring ground truth geometry information. The pre-trained depth model can be further fine-tuned for specific depth tasks and achieves superior results over training from scratch (see Fig. 1 right, where unsupervised pre-training leads to improved performance).
We conduct extensive experiments on the large-scale TartanAir [55], ScanNet [12] and RealEstate10K [75] datasets for depth estimation and Gaussian splatting tasks, as well as the recently introduced DL3DV [30] dataset, which features complex real-world scenes and thus is more challenging. Under various evaluation settings, our DepthSplat achieves state-of-the-art results. We also achieve promising feed-forward reconstruction results of large-scale or 360 scenes from up to 12 input views at $512\times 960$ resolutions. The strong performance on both tasks demonstrates the mutual benefits of connecting Gaussian splatting and depth.
@@ 2 Related Work
Feed-Forward Gaussian Splatting. Several feed-forward 3D Gaussian splatting models [7, 46, 8, 58, 48, 64, 72] have been proposed in literature thanks to its efficiency and ability to handle sparse views. In particular, pixelSplat [7] and Splatter Image [46] predict 3D Gaussians from image features, while MVSplat [8] encodes the feature matching information with cost volumes and achieves better geometry. However, it inherently suffers from the limitation of feature matching in challenging situations like texture-less regions and reflective surfaces. In this paper, we propose to integrate monocular features from pre-trained monocular depth models [66] for more robust depth prediction and 3D Gaussian reconstruction. Another line of work like LGM [48], GRM [64], and GS-LRM [72] relies significantly on the training data and compute, discarding explicit feature matching cues and learning priors purely from data. This makes them expensive to train (e.g., GS-LRM [72] is trained with 64 A100 GPUs for 2 days), while our model can be trained in 20 hours with 8 GPUs. Moreover, our Gaussian splatting module, in return, enables pre-training depth model from large-scale unlabeled datasets without the need for ground truth depths.
Depth and Gaussian Splatting. Flash3D [45] explores a pre-trained monocular depth model [36] for single-image 3D Gaussian reconstruction. In contrast, we target large-scale scene reconstruction from multiple views. TranSplat [71] leverages monocular depth to improve Gaussian reconstruction. However, a crucial difference is that TranSplat uses monocular features to refine the coarse depth predicted from a cost volume, which makes it inherently vulnerable to the error in the coarse stage and, thus, suffers from the error propagation issue. In contrast, we avoid this by combining cost volume and monocular features early, which also enables our method to perform significantly better than TranSplat (Tab. 5). There has been another line of work [10, 50] which applies an additional depth loss in the Gaussian splatting optimization process. We note that these two approaches (feed-forward vs. per-scene optimization) are orthogonal.
@@ 3 DepthSplat
Given $N$ input images $\{{\bm{I}}^{i}\}_{i=1}^{N},({\bm{I}}^{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ are the image sizes) with corresponding projection matrices $\{\mathbf{P}_{i}\}_{i=1}^{N}$, ($\mathbf{P}_{i}\in\mathbb{R}^{3\times 4}$, computed from the intrinsic and extrinsic matrices), our goal is to predict dense per-pixel depth $\mathbf{D}_{i}\in\mathbb{R}^{H\times W}$ and per-pixel Gaussian parameters $\{(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}_{j=1}^{H\times W\times N}$ for each image, where $\bm{\mu}_{j}$, $\alpha_{j}$, $\bm{\Sigma}_{j}$ and $\bm{c}_{j}$ are the 3D Gaussian’s position, opacity, covariance, and color information. As shown in Fig. 2, the core of our method is a multi-view depth model augmented with monocular depth features, where we obtain the position $\bm{\mu}_{j}$ of each Gaussian by unprojecting depth to 3D with camera parameters, and other Gaussian parameters are predicted by an additional lightweight head.
More specifically, our depth model consists of two branches: one for modeling feature matching information using cost volumes, and another for extracting monocular features from a pre-trained monocular depth network. The cost volumes and monocular features are concatenated together for subsequent depth regression with a 2D U-Net and a softmax layer. For the depth task, we train our depth model with ground truth depth supervision. Our full model for novel view synthesis is trained with the photometric rendering loss, which can also be used as an unsupervised pre-training stage for the depth model. In the following, we introduce the individual components.
@@ 3.1 Multi-View Feature Matching
Multi-View Feature Extraction. For $N$ input images, we first use a lightweight weight-sharing ResNet [25] architecture to get $s\times$ downsampled features for each image independently. To handle different image resolutions, we make the downsampling factor $s$ flexible by controlling the number of stride-2 $3\times 3$ convolutions. For example, the downsampling factor $s$ is 4 if two stride-2 convolutions are used and 8 if three are used. To exchange information across different views, we use a multi-view Swin Transformer [31, 60, 61] which contains six stacked self- and cross-attention layers to obtain multi-view-aware features $\{{\bm{F}_{i}}\}_{i=1}^{N}$, ${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C}$, where $C$ is the feature dimension. More specifically, cross-attention is performed between each reference view and other views. When more than two images ($N>2$) are given as input, we perform cross-attention between each reference view and its top-$2$ nearest neighboring views, which are selected based on their camera distances to the reference view. This makes the computation tractable while maintaining cross-view interactions.
@@ 3.3 Feature Fusion and Depth Regression
We also apply a hierarchical matching [23] architecture where an additional refinement step at 2×2$\times$2 × higher feature resolution is employed to improve the performance further. More specifically, based on the coarse depth prediction, we perform a correspondence search on the 2×2$\times$2 × higher feature maps within the neighbors of the 2×2$\times$2 × upsampled coarse depth prediction. Since we already have a coarse depth prediction, we only need to search a smaller range at the higher resolution, and thus, we construct a smaller cost volume compared to the coarse stage. Such a $2$-scale hierarchical architecture not only leads to improved efficiency since most computation is spent on low resolution, but also leads to better results thanks to the use of higher-resolution features [23]. Similar feature fusion and depth regression procedure is used to get higher-resolution depth predictions, which are subsequently upsampled [38] to the full resolution.
@@ 3.4 Gaussian Parameter Prediction
For the task of 3D Gaussian splatting, we directly unproject the per-pixel depth maps to 3D with the camera parameters as the Gaussian centers $\bm{\mu}_{j}$. To predict other remaining Gaussian parameters $\alpha_{j}$ (opacity), $\bm{\Sigma}_{j}$ (covariance) and $\bm{c}_{j}$ (color), we use an additional lightweight 2D U-Net [39], where it takes the concatenated image, depth and feature information as input and outputs all the remaining Gaussian parameters. With all the predicted 3D Gaussians, we can render novel views with the Gaussian splatting operation [27].
@@ 4 Experiments
Training Details. For depth experiments, we mainly follow the setup of UniMatch [61] for fair comparisons on the ScanNet [12] dataset. More specifically, we train our depth model on 4 GH200 GPUs for 100K iterations with a total batch size of 32. For the ablation experiments on depth task, we mainly use synthetic datasets (TartanAir [55] and VKITTI2 [5]) for training since they provide high-quality ground truth depths. For Gaussian splatting experiments, we consider both low-resolution and high-resolution images. In particular, for comparisons with previous methods [7, 8], we mainly use $256\times 256$ resolution RealEstate10K [75] and $256\times 448$ resolution DL3DV [30] datasets. We also report high-resolution ($512\times 960$) results on RealEstate10K and DL3DV datasets for qualitative evaluations. For experiments on the $256\times 256$ RealEstate10K dataset, we train our model on 8 GH200 GPUs for 150K iterations with a total batch size of 32, which takes about 20 hours. For experiments on the $256\times 448$ DL3DV [30] dataset, we evaluate on the official benchmark split with 140 scenes, while other remaining scenes are used for training. We fine-tune our RealEstate10K pre-trained model on 4 A100 GPUs for 100K iterations with a total batch size of 4, where the number of input views is randomly sampled from 2 to 6. We evaluate the model’s performance on different number of input views (2, 4, 6). For high-resolution results, we fine-tune our low-resolution pre-trained models on high-resolution images. More training details are provided in the appendix. Our code and training scripts are available at github.com/cvg/depthsplat to ease reproducibility.
Table 1: DepthSplat model variants. We evaluate different monocular backbones and different multi-view models ($1$-scale and 2-scale features for hierarchical matching as described in Sec. 3.3), where both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both depth and view synthesis tasks. Our model parameters, inference time, and memory consumption are provided in Tab. A.1 of the appendix.
Monocular Multi-View Depth (TartanAir) 3DGS (RealEstate10K)
@@
ViT-S $1$-scale 8.46 93.02 26.76 0.877 0.123
ViT-B $1$-scale 6.94 94.46 27.09 0.881 0.119
ViT-L $1$-scale 6.07 95.52 27.34 0.885 0.118
ViT-S $2$-scale 7.01 94.56 26.96 0.880 0.122
ViT-B $2$-scale 6.22 95.31 27.27 0.885 0.120
ViT-L $2$-scale 5.57 96.07 27.44 0.887 0.119
@@
full 8.46 93.02 26.76 0.877 0.123
w/o mono feature 12.25 88.00 26.27 0.866 0.130
w/o cost volume 11.34 90.02 23.09 0.761 0.187
single branch 11.26 90.84 25.99 0.858 0.134
ConvNeXt-T 10.50 91.13 26.28 0.867 0.130
Midas 9.53 91.61 26.40 0.869 0.129
DINO V2 8.93 92.49 26.68 0.874 0.125
Depth Anything V1 8.38 93.23 26.70 0.875 0.125
Depth Anything V2 8.46 93.02 26.76 0.877 0.123
@@ 4.1 Model Variants
We first study several different model variants for both depth estimation and view synthesis tasks. In particular, we explore different monocular backbones [66] (ViT-S, ViT-B, ViT-L) and different multi-view models ($1$-scale and $2$-scale). We conduct depth experiments on the large-scale TartanAir [55] synthetic dataset, which features both indoor and outdoor scenes and has perfect ground truth depth. The Gaussian splatting experiments are conducted on the standard RealEstate10K [75] dataset. Following community standards, we report the depth evaluation metrics [18] of Abs Rel (relative $\ell_{1}$ error) and $\delta_{1}$ (percentage of correctly estimated pixels within a threshold) and novel view synthesis metrics [27] of PSNR, SSIM, and LPIPS. The results in Tab. 1 demonstrate that both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both tasks.
From Tab. 1, we can also observe that better depth network architecture leads to improved view synthesis. In Tab. B.1 of the appendix, we conduct additional experiments to study the effect of different initializations for the depth network to the view synthesis performance. Our results show that a better depth model initialization also contributes to improved rendering quality. In summary, both better depth network architecture and better depth model initialization lead to improved novel view synthesis results.
@@ 4.2 Ablation and Analysis
Fusion Strategy. We compare with alternative strategies for fusing the monocular features for multi-view depth estimation. In particular, we compare with MVSFormer [6] which constructs the cost volume with monocular features. More specifically, we replace our multi-view feature extractor with a weight-sharing ViT model and use the ViT features to build the cost volume as done in MVSFormer. This leads to a single-branch architecture (single branch in Tab. 2), unlike our two-branch design where the monocular features are not used to build the cost volume. We can observe that our fusion strategy performs significantly better than the single-branch design, potentially because our two-branch design disentangles feature matching and obtaining monocular priors, which makes the learning task easier.
@@
DeMoN [51] 0.231 0.761 0.289
BA-Net [47] 0.161 0.346 0.214
DeepV2D [49] 0.057 0.168 0.080
NeuralRecon [44] 0.047 0.164 0.093
DRO [24] 0.053 0.168 0.081
UniMatch [61] 0.059 0.179 0.082
DepthSplat 0.044 0.119 0.059
@@
4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
By connecting Gaussian splatting and depth, our DepthSplat provides a way to pre-train the depth model in a fully unsupervised manner. In particular, we first train our full model on the large-scale unlabeled RealEstate10K dataset (contains $\sim$ 67K Youtube videos) with only the Gaussian splatting rendering loss (Eq. 2), without any direct supervision on the depth predictions. After pre-training, we take the pre-trained depth model and further fine-tune it to the depth task on the mixed TartanAir [55] and VKITTI2 [5] datasets with ground truth depth supervision. In Tab. 3, we evaluate the performance on both in-domain TartanAir test set and the zero-shot generalization on unseen ScanNet and KITTI datasets. We can observe that the benefit of pre-training is especially significant on the challenging datasets like TartanAir and KITTI. The results align well with the observation in [19] which suggests that the pre-training acts as regularization and guides the learning towards better minima and enables better generalization to out-of-distribution datasets. The visual comparison results are provided in Fig. C.1 of the appendix. Given the increasing popularity of view synthesis [74, 57] and multi-view generative models [42, 3], new multi-view datasets [30] and models [52] are gradually introduced, our approach provides a way to pre-train depth models on large-scale unlabeled multi-view image datasets.
@@ 4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
DepthSplat 27.44 0.887 0.119
Figure 5: Visual synthesis on RealEstate10K. Our DepthSplat performs significantly better than pixelSplat [7] and MVSplat [8] in challenging regions.
Figure 6: Depth results on unseen samples. Our depth model generalizes robustly to different scene types.

### v3

@@ Abstract
Gaussian splatting and single-view depth estimation are typically studied in isolation. In this paper, we present DepthSplat to connect Gaussian splatting and depth estimation and study their interactions. More specifically, we first contribute a robust multi-view depth model by leveraging pre-trained monocular depth features, leading to high-quality feed-forward 3D Gaussian splatting reconstructions. We also show that Gaussian splatting can serve as an unsupervised pre-training objective for learning powerful depth models from large-scale multi-view posed datasets. We validate the synergy between Gaussian splatting and depth estimation through extensive ablation and cross-task transfer experiments. Our DepthSplat achieves state-of-the-art performance on ScanNet, RealEstate10K and DL3DV datasets in terms of both depth estimation and novel view synthesis, demonstrating the mutual benefits of connecting both tasks. In addition, DepthSplat enables feed-forward reconstruction from 12 input views ($512\times 960$ resolutions) in 0.6 seconds.
@@ 1 Introduction
The integration of 3DGS with single-view depth estimation presents a compelling solution to overcome the individual limitations of each technique while at the same time enhancing their strengths. To this end, we introduce DepthSplat, which exploits the complementary nature of sparse-view feed-forward 3DGS and robust monocular depth estimation to improve the performance for both tasks.
@@
Thanks to our improved multi-view depth model, the quality of novel view synthesis with Gaussian splatting is also significantly enhanced (see Fig. 1 left). In addition, our Gaussian splatting module is fully differentiable, which requires only photometric supervision to optimize all model components. This provides a new, unsupervised way to pre-train depth prediction models on large-scale multi-view posed datasets without requiring ground truth geometry information. The pre-trained depth model can be further fine-tuned for specific depth tasks and achieves superior results over training from scratch (see Fig. 1 right, where unsupervised pre-training leads to improved performance).
We conduct extensive experiments on the large-scale TartanAir [58], ScanNet [14] and RealEstate10K [82] datasets for depth estimation and Gaussian splatting tasks, as well as the recently introduced DL3DV [32] dataset, which features complex real-world scenes and thus is more challenging. Under various evaluation settings, our DepthSplat achieves state-of-the-art results. We also achieve promising feed-forward reconstruction results of large-scale or $360^{\circ}$ scenes from 12 input views ($512\times 960$ resolutions) in 0.6 seconds. The strong performance on both tasks demonstrates the mutual benefits of connecting Gaussian splatting and depth.
@@ 2 Related Work
Feed-Forward Gaussian Splatting. Several feed-forward 3D Gaussian splatting models [7, 48, 9, 62, 51, 68, 77, 10, 8] have been proposed in literature thanks to its efficiency and ability to handle sparse views. In particular, pixelSplat [7] and Splatter Image [48] predict 3D Gaussians from image features, while MVSplat [9] encodes the feature matching information with cost volumes and achieves better geometry. However, it inherently suffers from the limitation of feature matching in challenging situations like texture-less regions and reflective surfaces. In this paper, we propose to integrate monocular features from pre-trained monocular depth models [70] for more robust depth prediction and 3D Gaussian reconstruction. Another line of work like LGM [51], GRM [68], and GS-LRM [77] relies significantly on the training data and compute, discarding explicit feature matching cues and learning priors purely from data. This makes them expensive to train (e.g., GS-LRM [77] is trained with 64 A100 GPUs for 2 days), while our model can be trained in 2 days with 4 GPUs. Moreover, our Gaussian splatting module enables pre-training depth model from large-scale multi-view posed datasets without ground truth depth supervision.
Depth and Gaussian Splatting. Flash3D [49] explores a pre-trained monocular depth model [39] for single-image 3D Gaussian reconstruction. In contrast, we target large-scale scene reconstruction from multiple views. TranSplat [76] leverages monocular depth to improve Gaussian reconstruction. However, a crucial difference is that TranSplat uses monocular features to refine the coarse depth predicted from a cost volume, which makes it inherently vulnerable to the error in the coarse stage and, thus, suffers from the error propagation issue. In contrast, we avoid this by combining cost volume and monocular features early, which performs significantly better than TranSplat (Tab. 6). Another line of work [12, 53] applies an additional depth loss to the Gaussian optimization process. We note that these two approaches (feed-forward vs. per-scene optimization) are orthogonal.
@@ 3 DepthSplat
Given $N$ input images $\{{\bm{I}}^{i}\}_{i=1}^{N},({\bm{I}}^{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ are the image sizes) with corresponding projection matrices $\{\mathbf{P}_{i}\}_{i=1}^{N}$, ($\mathbf{P}_{i}\in\mathbb{R}^{3\times 4}$, computed from the intrinsic and extrinsic matrices), our goal is to predict dense per-pixel depth $\mathbf{D}_{i}\in\mathbb{R}^{H\times W}$ and per-pixel Gaussian parameters $\{(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}_{j=1}^{H\times W\times N}$ for each image, where $\bm{\mu}_{j}$, $\alpha_{j}$, $\bm{\Sigma}_{j}$ and $\bm{c}_{j}$ are the 3D Gaussian’s position, opacity, covariance, and color. As shown in Fig. 2, key to our method is a multi-view depth model augmented with monocular depth features, where we obtain the position $\bm{\mu}_{j}$ of each Gaussian by unprojecting depth with camera parameters, and other Gaussian parameters are predicted by an additional head.
More specifically, our depth model consists of two branches: one for modeling feature matching using cost volumes, and another for extracting monocular features from a pre-trained monocular depth network. The cost volumes and monocular features are concatenated together for subsequent depth regression with a 2D U-Net and a softmax layer. For the depth task, we train our depth model with ground truth depth supervision. Our full model for novel view synthesis is trained with the rendering loss, which can also be used as an unsupervised pre-training stage for the depth model. We introduce the individual components below.
@@ 3.1 Multi-View Feature Matching
Multi-View Feature Extraction. For $N$ input images, we first use a lightweight weight-sharing ResNet [27] architecture to get $s\times$ downsampled features for each image independently. To handle different image resolutions, we make the downsampling factor $s$ flexible by controlling the number of stride-2 $3\times 3$ convolutions. For example, the downsampling factor $s$ is 4 if two stride-2 convolutions are used and 8 if three are used. To exchange information across different views, we use a multi-view Swin Transformer [34, 64, 65] which contains six stacked self- and cross-attention layers to obtain multi-view-aware features $\{{\bm{F}_{i}}\}_{i=1}^{N}$, ${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C}$, where $C$ is the feature dimension. When more than two images ($N>2$) are given as input, we perform cross-attention between each reference view and its top-$2$ nearest neighboring views, which are selected based on their camera distances to the reference view. This makes computation tractable with many input views.
@@ 3.3 Feature Fusion and Depth Regression
We also apply a hierarchical matching [25] architecture where an additional refinement step at 2×2$\times$2 × higher feature resolution is employed to improve the performance further. More specifically, based on the coarse depth prediction, we perform a correspondence search on the 2×2$\times$2 × higher feature maps within the neighbors of the 2×2$\times$2 × upsampled coarse depth prediction. Since we already have a coarse depth prediction, we only need to search a smaller range at the higher resolution, and thus, we construct a smaller cost volume compared to the coarse stage. Such a $2$-scale hierarchical architecture not only leads to improved efficiency since most computation is spent on low resolution, but also leads to better results thanks to the use of higher-resolution features [25]. Similar feature fusion and depth regression procedure is used to get higher-resolution depth predictions, which are subsequently upsampled with a DPT head [41] to the full resolution.
@@ 3.4 Gaussian Parameter Prediction
For the task of 3D Gaussian splatting, we directly unproject the per-pixel depth maps to 3D with the camera parameters as the Gaussian centers $\bm{\mu}_{j}$. To predict other remaining Gaussian parameters $\alpha_{j}$ (opacity), $\bm{\Sigma}_{j}$ (covariance) and $\bm{c}_{j}$ (color), we use an additional DPT head [41], where it takes the concatenated image, depth and feature information as input and outputs all the remaining Gaussian parameters. With all the predicted 3D Gaussians, we can render novel views with the Gaussian splatting operation [29].
@@ 4 Experiments
Training Details. For depth experiments on ScanNet [14], we train our depth model on 4×4$\times$4 × GH200 GPUs for 100K iterations with a total batch size of 32. For the ablation experiments on depth task, we mainly use synthetic datasets (TartanAir [58] and VKITTI2 [5]) for training since they provide high-quality ground truth depths. For Gaussian splatting experiments, we consider both low-resolution and high-resolution images. In particular, for comparisons with previous methods [7, 9], we mainly use $256\times 256$ resolution RealEstate10K [82] and $256\times 448$ resolution DL3DV [32] datasets. We also report high-resolution ($512\times 960$) results on RealEstate10K and DL3DV datasets for qualitative evaluations. For experiments on the $256\times 256$ RealEstate10K dataset, we train our model on 4×4$\times$4 × GH200 GPUs for 150K iterations with a total batch size of 32, which takes 1 day for the small model and 2 days for the large model. For experiments on the $256\times 448$ DL3DV [32] dataset, we fine-tune our RealEstate10K pre-trained model on 4×4$\times$4 × GH200 GPUs for 100K iterations with a total batch size of 4, where the number of input views is randomly sampled from 2 to 6. We evaluate the model’s performance on different number of input views (2, 4, 6). For high-resolution results, we fine-tune our low-resolution pre-trained models on high-resolution images. More training details are provided in the appendix. Our code and training scripts are available at github.com/cvg/depthsplat to ease reproducibility.
Table 1: DepthSplat model variants. We evaluate different monocular backbones and different multi-view models ($1$-scale and 2-scale features for hierarchical matching as described in Sec. 3.3), where both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both depth and view synthesis tasks. The inference time is measured on a single A100 GPU for two input views at $256\times 256$ resolutions.
Mono Multi-View Depth (TartanAir) 3DGS (RealEstate10K) Param (M) Time (s)
@@
ViT-S $1$-scale 8.46 93.02 26.84 0.878 0.122 38 0.050
ViT-B $1$-scale 6.94 94.46 27.06 0.882 0.119 115 0.053
ViT-L $1$-scale 6.07 95.52 27.23 0.885 0.116 347 0.061
ViT-S $2$-scale 7.01 94.56 27.12 0.884 0.119 41 0.062
ViT-B $2$-scale 6.22 95.31 27.34 0.887 0.116 120 0.070
ViT-L $2$-scale 5.57 96.07 27.47 0.889 0.114 354 0.079
@@
Full 8.46 93.02 26.84 0.878 0.122
w/o mono feature 12.25 88.00 26.04 0.864 0.134
w/o cost volume 11.34 90.02 23.24 0.766 0.184
ConvNeXt-T 10.50 91.13 26.11 0.865 0.134
Midas 9.53 91.61 26.46 0.872 0.128
DINO V2 8.93 92.49 26.76 0.877 0.123
Depth Anything V1 8.38 93.23 26.76 0.877 0.124
Depth Anything V2 8.46 93.02 26.84 0.878 0.122
@@ 4.1 Model Variants
We first study several different model variants for both depth estimation and view synthesis tasks. In particular, we explore different monocular backbones [70] (ViT-S, ViT-B, ViT-L) and different multi-view models ($1$-scale and $2$-scale). We conduct depth experiments on the large-scale TartanAir [58] synthetic dataset, which features both indoor and outdoor scenes and has perfect ground truth depth. The Gaussian splatting experiments are conducted on the standard RealEstate10K [82] dataset. Following community standards, we report the depth evaluation metrics [20] of Abs Rel (relative $\ell_{1}$ error) and $\delta_{1}$ (percentage of correctly estimated pixels within a threshold) and novel view synthesis metrics [29] of PSNR, SSIM, and LPIPS. The results in Tab. 1 demonstrate that both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both tasks. We name three different model sizes with “Small": “ViT-S monocular, 1-scale multi-view" , “Base": “ViT-B monocular, 2-scale multi-view", and “Large": “ViT-L monocular, 2-scale multi-view".
From Tab. 1, we can also observe that better depth network architecture leads to improved view synthesis. In the appendix, we conduct additional experiments to study the effect of different initializations for the depth network to the view synthesis performance. Our results show that a better depth model initialization also contributes to improved rendering quality. In summary, both better depth network architecture and better depth model initialization lead to improved novel view synthesis results.
@@ 4.2 Ablation and Analysis
Table 3: Monocular fusion strategy. Our simple concatenation performs surprisingly good compared to other alternatives.
Method Abs Rel $\downarrow$ $\delta_{1}\uparrow$
Cost volume from ViT 11.26 90.84
Explicit scale align 9.13 90.21
Attention fusion 8.40 92.95
Concatenation (Ours) 8.46 93.02
@@
DeMoN [54] 23.1 0.761 0.289
BA-Net [50] 16.1 0.346 0.214
DeepV2D [52] 5.7 0.168 0.080
NeuralRecon [47] 4.7 0.164 0.093
DRO [26] 5.3 0.168 0.081
UniMatch [65] 5.9 0.179 0.082
DepthSplat (w/o GS pre-train) 4.5 0.125 0.061
DepthSplat (w/ GS pre-train) 3.8 0.114 0.055
@@
Monocular Fusion Strategy. We compare with alternative strategies for fusing the monocular features for multi-view depth estimation in Tab. 3. In particular, we first compare with MVSFormer [6] that constructs the cost volume with monocular ViT features, which leads to a single-branch architecture. We can observe that our fusion strategy performs significantly better than the single-branch design, potentially because our two-branch design disentangles feature matching and monocular priors, which makes the learning task easier. We also compare with fusing the monocular depth maps (instead of monocular features in our model) by first aligning the monocular depth scale using the multi-view depth and then fusing the depth maps together with an additional network. Again, our simple concatenation performs better. We further compare our concatenation with the attention-based adaptive fusion and observe very similar results. Thus we opt to the simpler concatenation design.
@@
DepthSplat 27.47 0.889 0.114
Figure 5: Novel view synthesis on RealEstate10K. Our DepthSplat performs significantly better than pixelSplat [7] and MVSplat [9] in challenging regions.
4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
By connecting Gaussian splatting and depth, our DepthSplat provides a way to pre-train the depth model in a fully unsupervised manner. In particular, we first train our full model on the large-scale multi-view posed RealEstate10K dataset (contains $\sim$ 67K Youtube videos) with only the Gaussian splatting rendering loss (Eq. 2), without any direct supervision on the depth predictions. After pre-training, we take the pre-trained depth model and further fine-tune it to the depth task on the mixed TartanAir [58] and VKITTI2 [5] datasets with ground truth depth supervision. In Tab. 4, we evaluate the performance on both in-domain TartanAir test set and the zero-shot generalization on unseen ScanNet and KITTI datasets. We can observe that the benefit of pre-training is especially significant on the challenging datasets like TartanAir and KITTI. The results align well with the observation in [21] which suggests that the pre-training acts as regularization and guides the learning towards better minima and enables better generalization to out-of-distribution datasets. The visual comparison results are provided in the appendix. Given the increasing popularity of view synthesis [81, 61] and multi-view generative models [45, 3], new multi-view datasets [32] and models [55] are gradually introduced, our approach provides a way to pre-train depth models on large-scale multi-view posed image datasets.